# Predictively Oriented Posteriors

### v1

@@ Appendix C Proofs of Main results
$\pi_{\mathcal{B}_{r}}(\theta):=\frac{1(\theta\in\mathcal{B}_{r})\pi(\theta)}{\Pi(\mathcal{B}_{r})}.$
Now, the RHS of the inequality in Lemma 1 is true for any $Q\in\mathcal{P}(\Theta)$, and remains true for $Q=\Pi_{\mathcal{B}_{n}}$. Under this choice for $Q$, and for $\delta_{n}(Q^{k})=\frac{1}{n}\sum_{i=1}^{n}\int\delta(\theta_{1:k},x_{i})\mathsf{d}Q^{k}(\theta_{1:k})$, equation (33) in Lemma 1 can be restated as
$\mathbb{E}\left[\mathcal{D}_{S}(P_{Q_{n}},P_{0})\right]$ $\leq\mathbb{E}\left[\inf_{r>0}\left\{\mathop{\text{\large$\int$}}\nolimits L_{n}(\theta_{1:k})\mathsf{d}\Pi_{\mathcal{B}_{r}}^{k}(\theta_{1:k})-\mathcal{S}(P_{0},P_{0})+\frac{\text{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathbb{E}\left[\inf_{r>0}\left\{\mathop{\text{\large$\int$}}\nolimits S_{n}(\theta)\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta_{k})-\mathcal{S}(P_{0},P_{0})-\delta_{n}(\Pi_{\mathcal{B}_{r}}^{k})+\frac{\text{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{r>0}\left[\mathbb{E}\left\{\mathop{\text{\large$\int$}}\nolimits S_{n}(\theta)\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)-\mathcal{S}(P_{0},P_{0})-\delta_{n}(\Pi_{\mathcal{B}_{r}}^{k})+\frac{\text{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)-\mathbb{E}\delta_{n}(\Pi^{k}_{\mathcal{B}_{r}})+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n},$
where the last line follows since $\delta_{n}(\Pi^{k}_{\mathcal{B}_{r}})\geq 0$ for all $r>0$.
Note that $\text{KL}(Q^{k}\|\Pi^{k})=k\cdot\text{KL}(Q\|\Pi)$, and that, for $0\leq r\leq r_{0}$, over $\mathcal{B}_{r}$, $\mathcal{D}_{S}(P_{\theta},P_{0})\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r$, so that we obtain
$\mathbb{E}\left[\mathcal{D}_{S}(P_{Q_{n}},P_{0})\right]$ $\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n},$
$=\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}\|\Pi)}{\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\log\left\{1/\Pi(\mathcal{B}_{r})\right\}}{\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n},$
$\leq\inf_{r_{0}\geq r>0}\left\{\int\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{d_{S}\log(c_{S}/r)}{\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\inf_{r_{0}\geq r>0}\left\{r+\frac{d_{S}\log(c_{S}/r)}{\lambda_{n}}\right\}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
The remainder of the proof follows the same as in the proof of Lemma 5. ∎
@@
${\mathsf{JenGap}}:=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\mathcal{D}_{S}\left(P_{Q^{\star}},P_{0}\right)-{\mathsf{AppGap}}(Q^{\star},k)>0,$
with ${\mathsf{AppGap}}(Q^{\star},k)\geq 0$ under Assumptions 1-2. According to Assumption 6, there also exists some $\widetilde{Q}^{\star}$ such that $\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)<+\infty$ and
$\left|\int L(\theta_{1:k})\mathsf{d}Q^{\star}(\theta)-\int L(\theta_{1:k})\mathsf{d}\widetilde{Q}^{\star}(\theta)\right|\leq\frac{\mathsf{JenGap}}{4}\,.$
Denoting
${\mathsf{AppGap}}_{n}(Q,k):=\frac{1}{n}\left\{\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)-S\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)\right\}\,$,
we can re-arrange equation (33) in the proof of Lemma 1 (valid due to Assumptions 1-3) as
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$
@@
$\leq\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)-S\left(P_{0},x_{i}\right)\right)+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathbb{E}\left[\inf_{Q\in\mathcal{P}(\Theta)}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(S\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)-S\left(P_{0},x_{i}\right)\right)+{\mathsf{AppGap}}_{n}(Q,k)+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}\right]+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(S\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)-S\left(P_{0},x_{i}\right)\right)+{\mathsf{AppGap}}_{n}(Q,k)+\frac{\text{KL}(Q^{k}\|\Pi^{k})}{k\lambda_{n}}\right]\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
@@
$=\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}.$ (34)
Hence, since $\mathrm{KL}(Q^{k}\|\Pi^{k})=k\cdot\mathrm{KL}(Q\|\Pi)$, we can obtain:
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\inf_{Q\ll\Pi}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
@@
$\leq\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}\widetilde{Q}^{\star}(\theta),P_{0}\right)+{\mathsf{AppGap}}(\widetilde{Q}^{\star},k)+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\underbrace{\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}\widetilde{Q}^{\star}(\theta),P_{0}\right)+{\mathsf{AppGap}}(\widetilde{Q}^{\star},k)}_{\equiv\int L(\theta_{1:k})\mathsf{d}\widetilde{Q}^{\star}(\theta_{1:k})-S(P_{0},P_{0})}-\underbrace{\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)-{\mathsf{AppGap}}(Q^{\star},k)}_{\equiv\int L(\theta_{1:k})\mathsf{d}{Q}^{\star}(\theta_{1:k})-S(P_{0},P_{0})}$
$\quad+\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)+{\mathsf{AppGap}}({Q}^{\star},k)+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$\leq\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)+{\mathsf{AppGap}}({Q}^{\star},k)+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
@@
$=\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}\left(P_{Q^{\star}},P_{0}\right)+\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\mathcal{D}_{S}\left(P_{Q^{\star}},P_{0}\right)-{\mathsf{JenGap}}\right\}+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}-\frac{3}{4}\mathsf{JenGap}.$
Hence using Assumption 4, for $n$ large enough so that
$\frac{\textnormal{KL}(\widetilde{Q}^{\star}\|\Pi)}{\lambda_{n}}+\frac{k\lambda_{n}C_{k}^{2}}{n}\leq\frac{\mathsf{JenGap}}{4}\,$,
we have
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{\mathsf{JenGap}}{2}\,.$
Furthermore, from Lemma 5,
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]\leq C\frac{\log(n)}{n^{1/2}}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})$
so that for $n$ large enough
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]\geq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{\mathsf{JenGap}}{4}\,$,
so that for $n$ large enough but finite,
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{\mathsf{JenGap}}{2}$
@@
$\leq\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]-\frac{\mathsf{JenGap}}{4}$
@@
To obtain the second result, return to equation (34), and use the fact that $Q^{\star}\in\operatornamewithlimits{arginf}_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})$ satisfies $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$, to obtain
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+{\mathsf{AppGap}}(Q^{\star},k)+\frac{\text{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\frac{C}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}+{\mathsf{AppGap}}(Q^{\star},k).$
Choosing $\lambda_{n}\propto n^{1/2}/\log(n)$ then yields the stated result. ∎
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{Q\ll\Pi}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$=\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+{\mathsf{AppGap}}(Q^{\star},k)+\frac{\text{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq{\mathsf{AppGap}}(Q^{\star},k)+C\frac{\log(n)}{n^{1/2}},$
where the last inequality follows from the fact that $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$ and choosing $\lambda_{n}=(n/\log n)^{1/2}$. Since in this case $\mathcal{D}_{S}(P_{Q^{\star}},P_{0})=0$,
${\mathsf{JenGap}}=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-{\mathsf{AppGap}}(Q^{\star},k)>0,$
so that under Assumption 5, there exist an $n$ large enough so that
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$ $\leq{\mathsf{AppGap}}(Q^{\star},k)-\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+C\frac{\log(n)}{n^{1/2}}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})$
$=-{\mathsf{JenGap}}+C\frac{\log(n)}{n^{1/2}}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})$
$<-\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0}).$
Further, for $n$ large enough, from Lemma 5, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right]\geq-\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0}).$ Putting these inequalities together, we see that, for $n$ large enough,
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$ $<-\frac{\mathsf{JenGap}}{4}+\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})\leq\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right].$
@@
$\mathsf{d}\pi^{\star}_{r,j}(\theta)=\begin{cases}\frac{\pi(\theta)}{\Pi(\mathcal{B}_{r,j})}&\text{ if }\theta\in\mathcal{B}_{r,j}\\ 0&\text{ else }\end{cases}.$

### v2

@@ Appendix C Proofs of Main results
From the convexity in Assumption 1,
$\mathcal{D}_{S}(P_{Q},P_{0})\leq\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q(\theta).$
Applying bound in the RHS of Lemma 1 delivers
$\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}(P_{Q},P_{0})+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}Q(\theta)+\frac{\mathrm{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}.$ (28)
@@
$\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta):=\frac{1(\theta\in\mathcal{B}_{r})\mathsf{d}\Pi(\theta)}{\Pi(\mathcal{B}_{r})}.$
Now, the RHS of the inequality in (28) is true for any $Q\in\mathcal{P}(\Theta)$, and remains true for $Q=\Pi_{\mathcal{B}_{n}}$. For $0\leq r\leq r_{0}$, over $\mathcal{B}_{r}$, $\mathcal{D}_{S}(P_{\theta},P_{0})\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r$, so that we obtain
$\mathbb{E}\left[\mathcal{D}_{S}(P_{Q_{n}},P_{0})\right]$ $\leq\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}^{k}\|\Pi^{k})}{k\lambda_{n}}\right\}+\nu_{n},$
$=\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\mathrm{KL}(\Pi_{\mathcal{B}_{r}}\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$=\inf_{r>0}\left\{\int\mathcal{D}_{S}(P_{\theta},P_{0})\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{\log\left\{1/\Pi(\mathcal{B}_{r})\right\}}{\lambda_{n}}\right\}+\nu_{n},$
$\leq\inf_{r_{0}\geq r>0}\left\{\int\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+r\right\}\mathsf{d}\Pi_{\mathcal{B}_{r}}(\theta)+\frac{d_{S}\log(c_{S}/r)}{\lambda_{n}}\right\}+\nu_{n}$
$=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\inf_{r_{0}\geq r>0}\left\{r+\frac{d_{S}\log(c_{S}/r)}{\lambda_{n}}\right\}+\nu_{n}\,.$
The remainder of the proof follows the same as in the proof of Lemma 3.
A similar argument holds for $Q_{n}^{\dagger}$.
@@
${\mathsf{JenGap}}:=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\mathcal{D}_{S}\left(P_{Q^{\star}},P_{0}\right)>0.$
From Lemma 1, we can obtain:
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
@@
$=\inf_{Q\ll\Pi}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
@@
$\leq\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}(\theta),P_{0}\right)+\frac{\textnormal{KL}({Q}^{\star}\|\Pi)}{\lambda_{n}}+\nu_{n}.$ (29)
Adding and subtracting $\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})$ and using the definition of $\mathsf{JenGap}$, we can re-arrange (29) as
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\frac{\textnormal{KL}({Q}^{\star}\|\Pi)}{\lambda_{n}}+\nu_{n}-\mathsf{JenGap}.$
Since, by Assumption 4, $Q^{\star}$ is such that $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$, for $n$ large enough we have
$\frac{\textnormal{KL}({Q}^{\star}\|\Pi)}{\lambda_{n}}+\nu_{n}\leq\frac{1}{4}{\mathsf{JenGap}}\,$,
so that
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{3}{4}{\mathsf{JenGap}}\,.$
Furthermore, from the continuity of $\theta\mapsto P_{\theta}$ in Assumption 4, and the concentration in Lemma 3, we have for $n$ large enough but finite,
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]\geq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{\mathsf{JenGap}}{4}\,.$
Therefore, for $n$ large enough but finite,
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\frac{3}{4}{\mathsf{JenGap}}$
@@
$\leq\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q_{n}^{\dagger}(\theta),P_{0}\right)\right]-\frac{1}{2}{\mathsf{JenGap}}$
@@
To obtain the second result, return to equation (29), and use the fact that $Q^{\star}\in\operatornamewithlimits{arginf}_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})$ satisfies $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$, by Assumption 4, to obtain
$\mathbb{E}\left[\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q_{n}(\theta),P_{0}\right)\right]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+\frac{\text{KL}(Q^{\star}\|\Pi)}{\lambda_{n}}+\nu_{n}$
$\leq\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+\frac{C}{\lambda_{n}}+\nu_{n}.$
In the case of Assumption 2, $\nu_{n}=\log(n)/{n}^{1/2}$ so that choosing $\lambda_{n}\propto n^{1/2}/\log(n)$ yields the stated result. In the case of Assumption 2, $\nu_{n}=\min\{\log(n)/{n}^{1/2},r_{n}\}$, and again choosing $\lambda_{n}\propto{n}^{1/2}\log(n)$ also implies the stated result. ∎
@@
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]$ $\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{Q\ll\Pi}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$=\mathcal{D}_{S}(P_{Q^{\star}},P_{0})+\frac{C}{\lambda_{n}}+\nu_{n},$
where the equality follows from the fact that $\mathrm{KL}(Q^{\star}\|\Pi)<\infty$, and the last inequality follows by taking $\lambda_{n}\propto{n}^{1/2}/\log(n)$ so that $C/\lambda_{n}+\nu_{n}\leq C^{\prime}\cdot\nu_{n}$, for $n$ large enough. We remind the reader that $\mathcal{D}_{S}(P_{Q^{\star}},P_{0})=0$ under Assumption 5(i).
Since $\mathcal{D}_{S}(P_{Q^{\star}},P_{0})=0$ under convex recovery, we have
${\mathsf{JenGap}}=\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})>0,$
and there exist an $n$ large enough so that
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq C^{\prime}\nu_{n}\leq\frac{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})}{2}.$
Further, for $n$ large enough, as in the proof of Theorem 3, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right]>\frac{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})}{2}.$ Putting these inequalities together, we see that, for $n$ large enough,
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\frac{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})}{2}<\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}^{\dagger}},P_{0}\right)\right].$
@@
$\mathsf{d}\Pi^{\star}_{r,j}(\theta)=\begin{cases}\frac{\mathsf{d}\Pi(\theta)}{\Pi(\mathcal{B}_{r,j})}&\text{ if }\theta\in\mathcal{B}_{r,j}\\ 0&\text{ else }\end{cases}.$